OpenStudy (anonymous):

The gravitational force between two objects increases as mass A. decreases or distance decreases. B. decreases or distance increases. C. increases or distance decreases. D. increases or distance increases.

OpenStudy (ivancsc1996):

C.

OpenStudy (ivancsc1996):

\[\frac{ m _{1}m _{2}G }{ r ^{2} }=F _{g}\]This equation tells us that mass is directly proportional to the masses and inversly proportional to the distance between them. \[F _{g}\alpha m_{1}m _{2}\alpha r ^{-2}\]Therefore the answer must be C. increases or distance decreases